OpenStudy (adelyn_helps):

How many solutions does the system of equations have? A. The system has exactly one solution. B. The system has infinitely many solutions. C. The system has no solution.

OpenStudy (adelyn_helps):

would it be B?

OpenStudy (rebecca):

Nope. x=y x-y=2 x-y=0 x-y=0  i used elimination.